Equipment is purchased which has an initial cost of $180,000. It has a 10 year life and its salvage value is estimated to be $15,000. Determine the initial entry into the accounting equation. What would be added annually to the accounting equation? (not a value, just the name of the entry)
Answer #1

(a)

Initial entry will be:

DR Equipment $180,000

CR Cash $180,000

(Being equipment purchased for $180,000)

(b)

The annual addition will be the depreciation of the equipment, deducted from book value. If straight line method is assumed,

Annual depreciation ($) = (Cost - Salvage value) / Useful life = (180,000 - 15,000) / 10 = 165,000 / 10 = 16,500